OpenStudy (anonymous):

Find all real solutions of 1/3log2(x + 6) = log2^(3).

OpenStudy (alfie):

\[\huge \frac{1}{3}\log_2(x+6) = \log_23\] Like this?

OpenStudy (alfie):

\[\large \log_2(x+6)^{\frac{1}{3}} = \log_23 \rightarrow \sqrt[3]{x+6} =3\] Can you take it from here?

OpenStudy (anonymous):

no

OpenStudy (alfie):

x+6 = (3)^3

OpenStudy (alfie):

x+6 = 27... x = 21